how do you write an equation for this function
the output is 5 more than 2 times the input.find the input when the output is 17

Answer by Thanenjayan(7)

say the input is "x"(cuz input is variable)
so we will build up the equation
so the out put is 5 more than 2times the input,
therefore 2 times input, means 2x and 5 more, means 5 should be added with the 2 times input
therfore we get the equation 2x+5 = output.
Cheers we hav come up with the equation. So substitude this equation in the question.
2x+5= output
therefore 2x+5=17
2x= 17-5
x= 12/2
= 6